# Generative Active Adaptation for Drifting and Imbalanced Network Intrusion Detection

### v1

@@ 4. Evaluation
We evaluate NetGuard end-to-end performance (inclduing active learning and data augmentation) based on F1-score improvement under drift, FPR, FNR, Accuracy, and rare class F1 improvement. We also compare the performance scores and labeling cost of NetGuard with state-of-the-art concept drift adaptation. We also analyze whether augmenting rare classes enhances classification while measuring the fidelity of synthetic samples.
@@ 4.1. Datasets and Model Selection
To evaluate our framework, we explored two real-world network intrusion detection datasets: the CIC-IDS dataset (Sharafaldin et al., 2018) from the Canadian Institute of Cybersecurity and the UGR’16 dataset from an ISP in Spain (Maciá-Fernández et al., 2018). In our experiments, we designate one dataset as the source dataset (where training is performed) and the other as the target dataset (where testing is conducted). Both datasets are split into 70% training and 30% testing sets, which is a standard practice in machine learning. For CIC-IDS, we use CSV files with flow-based features to conduct two experimental setups (a) training on 70% of CIC-IDS 2017 (source) and testing on 30% of CIC-IDS 2018 (target), and (b) training on 70% of CIC-IDS 2018 (source) and testing on 30% of CIC-IDS 2017 (target). This bidirectional evaluation highlights concept drift and demonstrates our framework’s ability to adapt to shifting distributions. For UGR’16, we follow the same strategy using July network traffic as the source dataset for training and August as the target dataset for testing. This allows us to analyze temporal drift across the months of July and August in real-world ISP data. We employ a multi-layer perceptron (MLP) with two hidden layers (100 neurons each) (Florencio et al., 2018) to assess performance under evolving network conditions. For CIC-IDS, we subsample 100k records per year, noting poor performance in Web Attack, Infiltration, and FTP-BruteForce (2018) and Infiltration (2017) due to class imbalance (Section 5.2, Table 2). For UGR’16, we sample 100k records, where minority classes (Botnet and Spam) pose classification challenges. To generate high-quality synthetic samples, we use deep generative models: (a) TVAE (Tabular Variational Autoencoder) and (b) RealTabFormer (RTF) (Solatorio and Dupriez, 2023). TVAE leverages a variational autoencoder framework for structured tabular data, while RTF, a transformer-based model, employs sequence-to-sequence (Seq2Seq) modeling to capture feature dependencies and structured patterns in tabular data. The strength of the generated samples is set to be 3×3$\times$3 × and 10×10$\times$10 × the original count of that minority class.
@@ 4.2. Baselines
We compare NetGuard with the existing active learning approaches as well as other concept drift adaptation schemes such as CADE (Yang et al., 2021).For active learning, the first baseline is the most ubiquitously used Uncertainty sampling (Bilgic et al., 2010) that selects samples for which the model produces the most uncertain predictions (where the entropy computed from the softmax layer of a neural network is used for uncertainty prediction). The second baseline includes the Coreset (Sener and Savarese, 2018) approach that selects a subset of samples using K-means clustering approach by choosing each cluster’s center as a training point to ensure diversity and coverage in training. The third baseline is CLUE (Prabhu et al., 2021) that applies weighted-clustering in the feature space to select instances that are both highly uncertain and diverse to maximize informativeness. For each of the active learning methods, we set the analyst’s labeling budget to $0.1\%,\;0.5\%,\;\text{and }1\%$ of the testing dataset. Additionally, we present two constant baselines: (1) a scenario where no adaptation scheme is applied, representing a pure concept drift setting, and (2) a scenario where full adaptation is performed by training on 70% of the training dataset from the target domain. This evaluation helps determine how NetGuard performs when no retraining is conducted versus when the classifier is fully retrained using all available training data. In order to to gain better insights into the contribution of our end-to-end framework, we also compare with the state-of-the-art concept drift adaptation technique. Specifically, we utilize CADE (Yang et al., 2021) that originally trains a contrastive encoder to identify out-of-distribution (OOD) samples and assigns a OOD score to the samples. We adapt their concept drift adaptation scheme into our active learning setting in order select informative samples from the target dataset for retraining.
Table 3. NetGuard ’s end-to-end performance on CIC 2017/2018 datasets
Method Label Budget F1 (%) Accuracy (%) FPR FNR F1 (%) Accuracy (%) FPR FNR
No Adapt 0% 60.62 60.62 0.5495 0.3062 58.28 58.28 0.3803 0.4925
Coreset 0.1% 61.74 70.71 0.5453 0.2929 80.55 82.61 0.3734 0.1739
0.5% 66.49 72.31 0.3894 0.2769 84.06 85.79 0.3353 0.1421
1% 63.34 68.00 0.4427 0.3201 87.28 89.35 0.2860 0.1065
Uncertainty 0.1% 66.70 73.19 0.4424 0.2681 73.14 67.33 0.2142 0.3267
0.5% 67.73 76.13 0.4985 0.2387 82.25 80.98 0.2125 0.1902
1% 74.18 79.13 0.4009 0.2087 84.41 84.55 0.2716 0.1545
CLUE 0.1% 62.46 70.77 0.5292 0.2923 91.63 91.36 0.1406 0.0864
0.5% 75.24 77.71 0.3431 0.2230 94.38 93.90 0.0686 0.0610
1% 79.71 83.39 0.3443 0.1661 95.98 96.01 0.0564 0.0398
NetGuard AL (Ours) 0.1% 70.15 74.65 0.4359 0.2535 89.07 89.10 0.1992 0.1090
0.5% 77.17 80.08 0.3352 0.1991 94.23 94.04 0.1027 0.0596
1% 81.78 85.81 0.2686 0.1419 96.19 96.10 0.0499 0.0390
NetGuard (End-to-end) 1% 86.00 88.00 0.2450 0.1339 98.01 99.00 0.0500 0.0256
Full Adapt 100% 93.22 93.22 0.0904 0.0556 98.95 98.95 0.0107 0.0102
@@ 4.3.1. CIC 2017 and CIC 2018 Datasets
We evaluate NetGuard ’s performance in two key aspects: (1) the effectiveness of the active learning (AL) module, and (2) the end-to-end performance, which includes both AL and data augmentation using RTF.The results in Table 5 demonstrate the effectiveness of NetGuard AL and the end-to-end approach (NetGuard AL + Augmentation) across different labeling budgets and scenarios. NetGuard AL consistently outperforms other methods, achieving higher F1 scores and lower FPR and FNR. For example, with a 1% labeling budget, NetGuard AL achieves an F1 score of 81.78% for Train: 2017 → Test: 2018 and 96.19% for Train: 2018 → Test: 2017. The end-to-end approach further improves these results, reaching 86.00% and 98.01% F1 scores, respectively. This improvement is more pronounced with higher labeling budgets, as more samples enhance the quality of augmented data. It is worth mentioning that NetGuard AL also benefits from considering the distributional discrepancy between datasets, which is a key factor in its superior performance. In contrast, methods like Uncertainty and CLUE struggle due to their reliance on model performance for sample selection where the model is already over-fitted to the distorted and limited latent space leading to wrong estimation of informativeness for sample selection. For example, the Uncertainty model performs poorly because it tends to misclassify rare classes with low uncertainty (i.e. high confidence), leading to biased sampling towards majority classes. Similarly, CLUE’s dependency on model performance and assumption of static latent space as well as Coreset’s distance-based selection fail to address the distributional discrepancies effectively. In the scenario where training is on 2017 data and testing on 2018 data, the end-to-end approach achieves 86% F1 with a 1% labeling budget, close to the Full Adapt scenario that gives an upper bound of 93.22%. This is because the 2018 dataset is more challenging with more distributional shifts and diverse patters in contrast with the 2017 dataset. In the reverse scenario (Train: 2018 → Test: 2017), the end-to-end approach nearly matches the Full Adapt performance, highlighting its robustness in handling distributional shifts and achieving near-optimal results.
Table 4. Class-level F1-score with 3x data augmentation of rare classes for Train: 2017 $\rightarrow$ Test: 2018
Attack No Adapt 1% CLUE 1% NetGuard AL NetGuard Full Adapt
(End-to-end)
Benign 0.46 0.90 0.92 0.88 0.85
Bot 0.0 0.96 1.00 1.00 0
DDoS 0.0 0.73 0.86 0.92 0.99
DoS GoldenEye 0.39 0.91 0.96 0.96 0.99
DoS Hulk 0.90 0.89 0.96 0.98 0.99
DoS Slowhttptest 0.0 0.44 0.63 0.72 0.62
DoS slowloris 0.92 0.85 0.94 0.95 1.0
FTP-BruteForce 0.0 0.57 0.51 0.71 0.74
Infiltration 0.0 0.20 0.23 0.28 0.34
SSH-Bruteforce 0.90 1.00 0.99 1.00 1.0
Web Attack 0.0 0.24 0.43 0.50 0.8
Class-specific adaptation effectiveness.
@@
The results in Table 4 demonstrate the impact of NetGuard on class-level performance across different attack types. Since certain attack classes exhibited extremely low performance during the active learning stage of NetGuard , we applied augmentation specifically to those classes by generating three times the original strength using RTF.For Table 4 scenario those rare attack types include DoS Slowhttptest, FTP-BruteForce, Infiltration, and Web Attack. As shown, the F1-score for DoS Slowhttptest increased from 0.0 with no adaptation to 0.44 using CLUE and 0.63 with NetGuard AL, further improving to 0.72 with augmentation, exceeding Full Adapt (0.62) due to the added diversity, which enhanced classifier robustness. FTP-BruteForce saw substantial gains, rising from 0.0 to 0.51 with NetGuard AL and 0.71 with augmentation, closely matching Full Adapt (0.74). Infiltration, one of the most challenging classes due to its rarity, improved from 0.23 with NetGuard AL to 0.28 after augmentation but still trailed Full Adapt (0.34). Web Attack followed a similar trend. We emphasize a key observation: as we augment the dataset by generating additional samples for Web Attack, FTP-BruteForce, Infiltration, and DoS Slowhttptest and incorporating them into the original CIC-IDS2017 dataset for IDS training, the performance of most other classes remains consistently stable within a range of 0.0 to 0.20. In some cases, augmentation even improves the performance of non-augmented classes, such as DDoS, DoS Hulk, and DoS slowloris, demonstrating that the added diversity helps refine the classifier’s decision boundaries. This finding is significant as the augmentation does not degrade the overall IDS performance, thus validating the feasibility of targeted augmentation in enhancing underrepresented attack detection. We obtained similar improvement for the augmented class in reverse setting with Train: 2018 → Test: 2017 shown in Appendix.
@@
Another noteworthy observation is NetGuard effectiveness in handling novel attacks, particularly when dealing with novel attack variants. As seen in Figure 2, new DDoS attack variants emerged in the CIC-IDS2018 (DDoS UDP + DDoS HTTP) dataset compared to CIC-IDS2017 (with only DDoS HTTP). This trend is reflected in Table 4, where, under the Train: 2017 → Test: 2018 setting, DDoS detection without adaptation results in an F1-score of 0.0, indicating a complete failure to recognize the newly introduced attack patterns. However, with classifier’s performance significantly improves reaching an F1 score of 0.92 with NetGuard ’s end-to-end approach. These results highlight that NetGuard not only adapts to distribution shifts for similar attack types but is also highly effective in identifying novel attacks and new variants, making it a robust solution for real-world intrusion detection scenarios where emerging threats are a persistent challenge.
Comparison with OOD sample selection using CADE
As shown in Table 5, we compare NetGuard’s performance with CADE in the scenario where training is performed on 2017 and testing on 2018. To ensure a fair comparison, we adapt CADE’s contrastive learning approach as a sample selection criterion for active learning, where the OOD score of a test sample is computed based on its distance to the nearest known class, and samples with highest OOD scores are selected for active learning from the 2018 training dataset for retraining the classifier. It is observed that even with 40% of the samples, CADE + Retrain achieves an F1-score of 72.26%, whereas NetGuard end-to-end performance surpasses it with 81.78% F1 using only 1% labeled data. This is likely because of the concept drift between datasets as shown in Fig. 1, the distribution of the same attack class shifts in the new domain, breaking CADE’s assumption that samples from the same class remain similar. This shift can also lead to overlapping embeddings between attack and benign classes, further reducing CADE’s ability to distinguish between them.
4.3.2. UGR’16 Datasets for July and August
Overall adaptation effectiveness.

### v2

@@ 4. Evaluation
We evaluate NetGuard end-to-end performance (including prior selection and data augmentation) based on F1-score improvement under drift, FPR, FNR, Accuracy, and rare class F1 improvement. We also compare the performance scores and labeling cost of NetGuard with state-of-the-art concept drift adaptation. We also analyze whether augmenting rare classes enhances classification while measuring the fidelity of synthetic samples. For synthetic data augmentation, we conduct experiments on an NVIDIA A100 GPU in Google Cloud and benchmark its utilization and latency overhead. Due to space limitations, we provide the computational cost analysis of NetGuard modules in the Appendix Table 9.
@@ 4.1. Datasets and Model Selection
To evaluate NetGuard , we use two popular NIDS datasets— CIC-IDS (2017/2018) from Canadian Institute of Cybersecurity  (Sharafaldin et al., 2018) and real-world UGR’16 network flows from an ISP in Spain  (Maciá-Fernández et al., 2018) capturing temporal drift (time-based shifts) and spatial drift (cross-domain/infrastructure differences). For CIC-IDS, we test bidirectional scenarios (1) Train on 2017 → Test on 2018 capturing both temporal and spatial drifts  [^1] CIC 2017 uses small enterprise LAN with 4 attack, 10 victim machines vs. CIC 2018’s large hybrid LAN/cloud setup with 10 attack, ¿ 400 victim machines, NAT/DHCP), and the reverse scenario (2) Train on 2018 → Test on 2017 that mainly reflects spatial drift despite the reversed temporal order. It is worth mentioning that CIC-IDS is commonly used for studying network drift and imbalance (Amalapuram et al., 2024; Yang et al., 2021; Chen et al., 2023c), and similarly, UGR’16 serves as a benchmark for distribution shift and traffic generation  (Yin et al., 2022; Xu et al., 2024; Zoppi et al., 2021), making these well-suited for NetGuard . Adversarial drift is out of scope, as generating protocol-compliant attack traffic requires specialized testbeds (Sheatsley et al., 2021). We use a standard 70/30 train-test split, with the training set as source domain and test set as target domain to evaluate drift. We employ an MLP (two 100-neuron hidden layers) for CIC-IDS to model non-linear traffic patterns  (Wei et al., 2023; Nadeem et al., 2023; Amalapuram et al., 2022b; Florencio et al., 2018) and XGBoost for UGR’16  (Bhandari et al., 2020; Attia et al., 2020b; Gao et al., 2024) to handle class imbalance, using 100k subsamples per dataset with Web Attack, Infiltration, and FTP-BruteForce (CIC-IDS) and Botnet/Spam (UGR’16) as minority classes. Our model choices align with recent comparisons of neural and non-neural classifiers in prior cybersecurity applications  (Chen et al., 2023c). To To address imbalance, we generate high-quality synthetic samples at 3×/10× ratios using deep generative models: (a) TVAE (Tabular Variational Autoencoder) (Apellaniz et al., 2024) and (b) RealTabFormer (RTF) (Solatorio and Dupriez, 2023). TVAE leverages a variational autoencoder framework for structured tabular data, while RTF, a transformer-based model, employs sequence-to-sequence (Seq2Seq) modeling to capture feature dependencies and structured patterns in tabular data.
Table 3. MLP performance on CIC 2017/2018 datasets
Method Label F1(%) FNR FPR Acc.(%) F1(%) FNR FPR Acc.(%)
None 0% 60.62 0.3062 0.5495 60.62 58.28 0.4925 0.3803 58.28
Coreset 0.1% 61.74 0.2929 0.5453 70.71 80.55 0.1739 0.3734 82.61
0.5% 66.49 0.2769 0.3894 72.31 84.06 0.1421 0.3353 85.79
1% 63.34 0.3201 0.4427 68.0 87.28 0.1065 0.2860 89.35
Uncert. 0.1% 66.70 0.2681 0.4424 73.19 73.14 0.3267 0.2142 67.33
0.5% 67.73 0.2387 0.4985 76.13 82.25 0.1902 0.2125 80.98
1% 74.18 0.2087 0.4009 79.13 84.41 0.1545 0.2716 84.55
CLUE 0.1% 62.46 0.2923 0.5292 70.77 91.63 0.0864 0.1406 91.6
0.5% 75.24 0.2230 0.3431 77.71 94.38 0.0610 0.0686 93.90
1% 79.71 0.1661 0.3443 83.39 95.98 0.0398 0.0564 96.01
CADE+ Retrain 0.1% 33.03 0.6278 0.0659 43.22 35.38 0.5875 0.0546 49.10
0.5% 24.95 0.7128 0.0723 37.95 32.0 0.5754 0.0580 49.38
1% 39.36 0.5823 0.0632 46.00 40.0 0.5014 0.0534 53.44
ReCDA+ Retrain 0.1% 32.81 0.5669 0.1583 4.30 44.51 0.4878 0.2122 52.62
0.5% 33.86 0.5480 0.1435 45.21 45.94 0.4378 0.1447 51.22
1% 34.96 0.5447 0.1411 45.53 54.23 0.4270 0.1692 57.129
NetGuard PS 0.1% 70.15 0.2535 0.4359 74.65 89.07 0.1090 0.1992 89.10
0.5% 77.17 0.1991 0.3352 80.08 94.23 0.0596 0.1027 94.04
1% 81.78 0.1419 0.2686 85.81 96.19 0.0390 0.0499 96.10
NetGuard E2E 1% 86.00 0.1339 0.2450 88.00 98.01 0.0256 0.0500 99.00
Full 100% 93.22 0.0556 0.0904 93.22 98.95 0.0102 0.0107 98.95
@@ 4.2. Baselines
[…]
@@ 4.3.1. CIC 2017 and 2018 Datasets:
Overall adaptation effectiveness. We evaluate NetGuard ’s performance in two key aspects: (1) the effectiveness of the prior selection (PS) module, and (2) the end-to-end performance, which includes both PS and data augmentation using RTF. The results in Table 5 demonstrate the effectiveness of NetGuard PS and the end-to-end approach (NetGuard PS + Augmentation) across different labeling budgets and scenarios. NetGuard PS consistently outperforms other methods, achieving higher F1 scores and lower FPR and FNR. For example, with a 1% labeling budget, NetGuard PS achieves an F1 score of 81.78% for Train: 2017 → Test: 2018 and 96.19% for Train: 2018 → Test: 2017. Sensitivity experiments on GMM parameters for NetGuard PS, along with PCA visualizations illustrating its sampling strategy for minority classes, are provided in the Appendix. The end-to-end approach further improves these results, reaching 86.00% and 98.01% F1 scores, respectively. This improvement is more pronounced with higher labeling budgets, as more samples enhance the quality of augmented data. It is worth mentioning that NetGuard PS also benefits from considering the distributional discrepancy between datasets, which is a key factor in its superior performance. In contrast, methods like Uncertainty and CLUE struggle due to their reliance on model performance for sample selection, where the model is already over-fitted to the distorted and limited latent space, leading to the wrong estimation of informativeness for sample selection. For example, the Uncertainty model performs poorly because it tends to misclassify rare classes with low uncertainty (i.e., high confidence), leading to biased sampling towards majority classes. Similarly, CLUE’s dependency on model performance and assumption of static latent space, as well as Coreset’s distance-based selection, fail to address the distributional discrepancies effectively. In the scenario where training is on 2017 data and testing on 2018 data, the end-to-end approach achieves 86% F1 with a 1% labeling budget, close to the Full Adapt scenario that gives an upper bound of 93.22%. This is because the 2018 dataset is more challenging with more distributional shifts and diverse patterns in contrast with the 2017 dataset. In the reverse scenario (Train: 2018 → Test: 2017), the end-to-end approach nearly matches the Full Adapt performance, highlighting its robustness in handling distributional shifts and achieving near-optimal results.
@@
Class-specific adaptation effectiveness. The results in Table 4 demonstrate the impact of NetGuard on class-level performance across different attack types. Since certain attack classes exhibited extremely low performance during the prior selection stage of NetGuard , we applied augmentation specifically to those classes by generating three times the original strength using RTF. For Table 4 scenario, those rare attack types include DoS Slowhttptest, FTP-BruteForce, Infiltration, and Web Attack. As shown, the F1-score for DoS Slowhttptest increased from 0.0 with no adaptation to 0.44 using CLUE and 0.63 with NetGuard PS, further improving to 0.72 with augmentation, exceeding Full Adapt (0.62) due to the added diversity, which enhanced classifier robustness. FTP-BruteForce saw substantial gains, rising from 0.0 to 0.51 with NetGuard PS and 0.71 with augmentation, closely matching Full Adapt (0.74). Infiltration, one of the most challenging classes due to its rarity, improved from 0.23 with NetGuard PS to 0.28 after augmentation but still trailed Full Adapt (0.34). Web Attack followed a similar trend. We emphasize a key observation: as we augment the dataset by generating additional samples for Web Attack, FTP-BruteForce, Infiltration, and DoS Slowhttptest and incorporating them into the original CIC-IDS2017 dataset for IDS training, the performance of most other classes remains consistently stable within a range of 0.0 to 0.20. In some cases, augmentation even improves the performance of non-augmented classes, such as DDoS, DoS Hulk, and DoS slowloris, demonstrating that the added diversity helps refine the classifier’s decision boundaries. This finding is significant as the augmentation does not degrade the overall IDS performance, thus validating the feasibility of targeted augmentation in enhancing underrepresented attack detection. We obtained similar improvement for the augmented class in the reverse setting with Train: 2018 → Test: 2017, shown in the Appendix 7
Table 4. NetGuard Class-wise E2E results (Train:2017 $\rightarrow$ Test:2018)
F1 scores (in %) for 1% labeling cost
Attack Class No Adapt CLUE CADE ReCDA NetGuard Full Adapt
Benign 0.46 0.90 0.47 0.47 0.88 0.85
Bot 0.0 0.96 0.0 0.0 1.00 0
DDoS 0.0 0.73 0.36 0.0 0.92 0.99
DoS GoldenEye 0.39 0.91 0.51 0.41 0.96 0.99
DoS Hulk 0.90 0.89 0.93 0.94 0.98 0.99
DoS Slowhttptest 0.0 0.44 0.00 0.00 0.72 0.62
DoS slowloris 0.92 0.85 0.97 0.95 0.95 1.0
FTP-BruteForce 0.0 0.57 0.00 0.00 0.71 0.74
Infiltration 0.0 0.20 0.003 0.15 0.28 0.34
SSH-Bruteforce 0.90 1.00 0.77 0.89 1.00 1.0
Web Attack 0.0 0.24 0.31 0.07 0.50 0.8
@@
.Another key observation is NetGuard effectiveness in handling novel attacks, particularly when dealing with novel attack variants. As seen in Figure 2, new DDoS attack variants emerged in the CIC-IDS2018 (DDoS UDP + DDoS HTTP) dataset compared to CIC-IDS2017 (with only DDoS HTTP). This trend is reflected in Table 4, where, under the Train: 2017 → Test: 2018 setting, DDoS detection without adaptation results in an F1-score of 0.0, indicating a complete failure to recognize the newly introduced attack patterns. However, with classifier’s performance significantly improves reaching an F1 score of 0.92 with NetGuard ’s end-to-end approach. These results highlight that NetGuard not only adapts to distribution shifts for similar attack types but is also highly effective in identifying novel attacks and new variants, making it a robust solution for real-world intrusion detection scenarios where emerging threats are a challenge.
Comparison with OOD sample selection using CADE and ReCDA. To ensure a fair baseline, we adapt CADE (Yang et al., 2021) and ReCDA  (Yang et al., 2024) as sample selection strategies, denoted as CADE+Retrain and ReCDA+ Retrain. Both select target-domain samples based on drift-aware scoring within a fixed labeling budget, combine them with full source data, and retrain an MLP classifier. CADE+Retrain uses a contrastive autoencoder to identify drifted samples via Mahalanobis distance in latent space. It retains the original encoder but retrains the classifier on raw features. Default parameters include: $\tau=0.5$, $\sigma=0.7$, $\text{TMAD}=2$, and $b=1.4826$. ReCDA+Retrain uses unsupervised scoring by ranking cosine similarity ratios in original vs. embedding space to K=2K{=}2 cluster centroids. A shallow encoder trained on source data guides selection. Parameters include $\tau=0.5$, $\sigma=0.7$, $\delta$ (selection ratio), and $\epsilon=10^{-5}$. As shown in Table 3, NetGuard outperforms both (81.78% F1 vs. 81.35% for CADE and 72.47% for ReCDA) using only 1% labeled samples from 100k data points. This improvement arises from significant concept drift between the source and target domains, illustrated in Fig. 1. As illustrated, attack class distributions shift substantially across domains, and embeddings of benign and attack classes begin to overlap in the target domain. These changes violate the core assumptions behind existing methods: CADE assumes intra-class embedding similarity, and ReCDA assumes embedding consistency. CADE achieves low FPR due to contrastive learning enforcing a closed-world boundary—drifted attacks near benign clusters are misclassified as benign. This reduces recall and F1 under drift. NetGuard handles open-world drift by generating synthetic samples that preserve overlap between classes while maintaining discriminative features. At 1% labeling, NetGuard achieves superior AUROC scores (0.91/0.99) compared to CADE+Retrain (0.68/0.81) and ReCDA+Retrain (0.67/0.78) for 2017→2018 and 2018→2017 transfers, showing NetGuard’s stronger ability to differentiate between multiple classes—unlike CADE and ReCDA, which lack full multi-class support (CADE experiments are limited to 4 classes in CIC-IDS and ReCDA supports only binary classes). We also experimented with CADE using a relaxed labeling strategy by selecting all drifted samples with parameter TMAD ¿ 2 (22,085 samples,  22% labeling cost). Despite this higher labeling cost, CADE achieved only 56.36% accuracy, 0.794 AUROC, 0.0503 FPR, and 0.4606 FNR, still underperforming compared to NetGuard with just 1% labeled data. Prior work shows that false positives increase analyst workload (Sommer and Paxson, 2010), and labeling capacity is limited to 80 samples/day (Braun et al., 2024). By using just 1% labeled data, NetGuard reduces annotation effort with high accuracy.
4.3.2. Improved interpretability with NetGuard via Trustee
To interpret the functionality of NetGuard , we used an interpretable rule-based module, Trustee (Jacobs et al., 2022), which extracts human-readable decision rules by approximating a trained classifier’s predictions using compact decision sets. We applied Trustee to analyze rules generated in the CIC-IDS setting (Train on 2017, Test on 2018) under two adaptation strategies: (1) NetGuard PS with only 1% labeled samples from the target domain, and (2) Full Adaptation using 100% of the target data. Our analysis reveals a strong alignment between the two sets of rules. In both cases, key security-relevant features such as PSH Flag Count, Active Mean, and Fwd IAT Mean consistently appear in the decision logic for identifying attacks like Infiltration and Web Attack. These overlaps confirm that NetGuard ’s prior selection (PS) effectively identifies representative and discriminative samples even with minimal supervision. By emphasizing flow-level and timing-based features, NetGuard maintains interpretability while achieving high detection performance. A detailed comparison of Trustee-derived rules across No Adaptation, NetGuard PS, and Full Adaptation is provided in the Appendix (Table 10).
4.3.3. UGR16: